# On the self-consistency of compact objects in Lorentz-violating gravity theories

### v1

@@ Abstract
Self-consistent solutions in Lorentz-violating gravity theories require the simultaneous satisfaction of: (i) the corresponding Einstein field equations, (ii) the matter field equations, and (iii) the Lorentz-violating field equations. In vacuum states, the dynamics of Lorentz-violating tensor fields may reduce to geometric constraints, potentially precluding entire classes of compact objects. These constraints are crucial for ensuring physical consistency in Lorentz-violating frameworks, as they eliminate metric families incompatible with the anisotropies induced by spontaneous Lorentz symmetry breaking. We investigate the criteria governing the emergence of these geometric constraints and analyze their consequences. Our analysis establishes a consistency framework for evaluating compact objects in these theories, demonstrating that several previously reported solutions in Lorentz-violating gravity models are physically inadmissible.
@@
Introduction. General relativity (GR) remains the most successful theory for describing gravitational interaction at classical scales, grounded in the symmetries of diffeomorphism and local Lorentz invariance. However, the unification of GR with Quantum Mechanics still represents one of the greatest unsolved challenges in theoretical physics. This motivates the plausible conjecture that modifications to the foundations of GR may pave the way toward a unified theory. The typical energy scale at which such a unified framework should becomes relevant is at the so-called Planck scale ($M_{\text{Pl}}=\sqrt{\hbar c/G}\approx 1.22\times 10^{19}\text{ GeV}/c^{2}$), far exceeding any observational regime accessible today. Nevertheless, theoretical extensions of GR predict the existence of low-energy “relic signatures” [1, 2].
@@
These signatures, which might arise from quantum gravitational effects, could manifest as subtle deviations either in experiments of the Standard Model or GR. In this context, the possibility of Lorentz symmetry violation in an underlying theory—such as string theory [3, 4]—becomes particularly relevant, especially in the search for observational signatures of new physics. These questions can be investigated independently of specific models through effective field theory approaches [5]. Thus, experimental tests of Lorentz violation in gravity not only validate extensions of GR but may also reveal evidence of physics beyond the current paradigm [6, 7].
@@
A promising approach to investigate Lorentz violation in pure gravity involves analyzing anisotropy effects induced in compact objects by tensor fields that acquire a preferred spacetime direction—a phenomenon arising from spontaneous Lorentz symmetry breaking [8]. In this mechanism, a potential term drives the dynamic evolution of the tensor field, leading to a nonzero vacuum expectation value (VEV). While being spontaneous, this violation preserves the Bianchi identities, despite implying both local Lorentz and diffeomorphism violation [5]. This compatibility with the geometric structure of GR makes spontaneous Lorentz symmetry breaking a natural candidate for exploring violation effects in strong gravity regimes, such as those found in neutron stars or black holes.
@@
The aim of this letter is to stress out a point, sometimes missing in the discussion on compact objects in Lorentz-violating gravity models, namely the self-consistency of compact objects in these frameworks. Crucially, self-consistent solutions of Lorentz-violating scenarios must simultaneously satisfy: (i) the corresponding Einstein field equations, (ii) the matter field equations, and (iii) the Lorentz-violating field equations. Remarkably, under certain conditions, the latter in a vacuum state can reduce to geometric constraints governing the metrics of compact objects. These constraints impose stringent requirements that rule out the emergence of families of compact objects and invalidate some previously reported solutions of Lorentz-violating gravity models.
@@
In the following, we analyze the emergence and implications of geometric constraints induced by spontaneous Lorentz violation in two prominent frameworks: (i) via a vector field (the bumblebee field) [5, 8], and (ii) through an antisymmetric rank-2 tensor field (sometimes known as the Kalb-Ramond field) [9]. The non-zero VEV of these fields introduces spacetime anisotropies, leading to small deviations in compact object solutions—including black holes [10, 11, 12, 13], neutron stars [14], and wormholes [15]—compared to standard GR predictions. Our analysis plays a crucial role in the search for viable compact objects within Lorentz-violating gravity, as it allows us to probe various families of metric functions (including those mentioned above) and determine which are permitted by the anisotropies induced by Lorentz-violating fields.
@@
Lorentz-violating framework. The action that encodes Lorentz-violating effects from a vector field (bumblebee field) [5] and from an antisymmetric rank-2 tensor [9] in Riemann-spacetime limit can be written as
@@
where $\mathfrak{B}_{\mu\nu}=\partial_{\mu}\mathfrak{B}_{\nu}-\partial_{\nu}\mathfrak{B}_{\mu}$ and $H_{\lambda\mu\nu}=\partial_{\lambda}B_{\mu\nu}+\partial_{\mu}B_{\nu\lambda}+\partial_{\nu}B_{\lambda\mu}$ are, respectively, the field strength of the bumblebee and of the antisymmetric rank-2 tensor, $\tilde{V}(Y)$ and $V(X)$ are self-interaction potentials that drive the breaking of the Lorentz symmetry with $Y=\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}\pm\mathfrak{b}^{2}$, $X=B_{\mu\nu}B^{\mu\nu}\pm b^{2}$, where $\mathfrak{b}^{2}$ and $b^{2}$ are positive real constants and the $\pm$ sign denotes whether the nature of the vector $\mathfrak{B}^{\mu}$ and of the antisymmetric tensor ${B}_{\mu\nu}$ is timelike or spacelike. Furthermore, the additional terms in the above Lagrangians account for all possible non-minimal couplings between the Lorentz-violating fields and spacetime curvature, with $R_{\alpha\beta\mu\nu}$ and $R_{\alpha\beta}$ denoting, respectively the Riemann and Ricci tensors and $\tilde{\xi}_{1,2}$ and ${\xi}_{1,2,3}$ are coupling constants. We recall that the bumblebee field does not couple to the Riemann tensor, as established in Ref. [5]. Moreover, we do not consider any coupling between the Lorentz-violating fields and the matter sector.
We assume that the self-interaction potentials $V$ and $\tilde{V}$ drive the formation of a nonzero vacuum value
$\langle\mathfrak{B}_{\mu}\rangle=\mathfrak{b}_{\mu},\quad\langle B_{\mu\nu}\rangle=b_{\mu\nu},$ (4)
which breaks local Lorentz and diffeomorphism symmetry. The minimum condition, i.e., $\langle V\rangle=0$ and $\langle\tilde{V}\rangle=0$ , is guaranteed when
$\mp\mathfrak{b}^{2}\equiv g^{\mu\nu}\mathfrak{b}_{\mu}\mathfrak{b}_{\nu},\quad\mp b^{2}\equiv g^{\mu\alpha}g^{\nu\beta}b_{\mu\nu}b_{\alpha\beta},$ (5)
where $\langle\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}\rangle=\mathfrak{b}_{\mu}\mathfrak{b}^{\mu}$ and $\langle B_{\mu\nu}B^{\mu\nu}\rangle=b_{\mu\nu}b^{\mu\nu}$. It is worth emphasizing that in the vacuum state, the terms coupling the Lorentz-violating fields to the curvature scalar (On the self-consistency of compact objects in Lorentz-violating gravity theories)-(On the self-consistency of compact objects in Lorentz-violating gravity theories) can be disregarded, as they can be absorbed into the Einstein-Hilbert term through a rescaling of the gravitational constant, i.e., $\kappa\to\kappa(1\pm\tilde{\xi}_{1}\mathfrak{b}^{2}\pm\xi_{1}b^{2})^{-1}$. Since the focus of this Letter is on the solutions influenced by the anisotropies induced by the VEV of the Lorentz-violating fields, henceforth $\xi_{1}=\tilde{\xi}_{1}=0$.
@@
where $T_{\mu\nu}=T_{\mu\nu}^{M}+T_{\mu\nu}^{LV}$ is the (total) energy-momentum tensor, with $T_{\mu\nu}^{M}$ and $T_{\mu\nu}^{LV}$ denoting the energy-momentum tensor of the matter fields and of the Lorentz-violating fields, respectively. We remark that $T_{\mu\nu}^{LV}$ is constructed by varying $S_{LV}=\int d^{4}x\sqrt{-g}(\mathcal{L}_{\mathfrak{B}}+\mathcal{L}_{{B}})$, containing, therefore, all the terms depending on $\mathfrak{B}_{\mu}$, $B_{\mu\nu}$ and their derivatives, along with all the non-minimal curvature-couplings controlled by $\tilde{\xi}_{2}$ and $\xi_{2,3}$. Additionally, by varying the action (1) with respect to the matter fields (collectively denoted by $\psi_{m}$), one obtains the equations of motion of the matter fields. Independently, by varying the action (1) with with respect to the Lorentz-violating fields, namely $\mathfrak{B}_{\mu}$ and $B_{\mu\nu}$, one obtains the field equations for the bumblebee field and for the antisymmetric rank-2 tensor field, respectively given by
$\nabla_{\alpha}\mathfrak{B}^{\alpha\nu}=$  $\,2\tilde{V}_{Y}\mathfrak{B}^{\nu}-\frac{\tilde{\xi}_{2}}{\kappa}\mathfrak{B}_{\mu}R^{\mu\nu},$ (7)
$\nabla_{\alpha}H^{\alpha\mu\nu}=$  $\,4V_{X}B^{\mu\nu}+\frac{2\xi_{2}}{\kappa}B_{\alpha}{}^{[\mu}R^{\nu]\alpha}-\frac{2\xi_{3}}{\kappa}B_{\alpha\beta}R^{\alpha\beta\mu\nu}.$ (8)
@@
We remark that, these consistency conditions are equivalently enforced by the conservation of the total energy-momentum tensor. By taking the covariant divergence of the Einstein equations and using the contracted Bianchi identity, $\nabla_{\mu}G^{\mu\nu}=0$, one obtains an on-shell condition that must hold:
$\nabla^{\mu}(T_{\mu\nu}^{M}+T_{\mu\nu}^{LV})=0.$ (9)
Once the equations of motion for matter and Lorentz-violating fields are satisfied, and using appropriate Bianchi identities, the above condition holds. Consequently, energy-momentum tensor conservation is guaranteed [5, 9].
Compact objects in Lorentz-violating gravity. The search for astrophysical objects in Lorentz-violating scenarios is typically based on the assumption that the Lorentz-violating fields are frozen at their VEVs. Depending on the assumption about these VEVs, different families of compact objects may arise in the Lorentz-violating scenario. It is important to point out that, imposing a particular form for the VEVs can turn the above dynamical equations into geometrical constraints, ruling out families of compact objects. These geometrical constraints emerge whenever the left-hand side of the eqs. (7)-(8) vanish. This can happen either for vanishing or (covariant) divergence-free field strength configurations. The resulting equations, for non-vanishing VEVs, imposes constraints on the Ricci and Riemann tensor.
Notably, several compact objects reported as solutions of Lorentz-violating models are based on the former assumption, namely the field strengths of the bumblebee and of the antisymmetric rank-2 tensor, in the vacuum state,—respectively $\mathfrak{b}_{\mu\nu}$ and $h_{\alpha\beta\gamma}$—, vanish. In the following we will show that this choice can, instead of allowing, forbid the emergence of some of these compact objects. We will additionally see that geometric constraints are stringent, and, for particular choices of VEV states, only a few classes of metrics are allowed.

### v2

@@ Abstract
Self-consistent solutions in Lorentz-violating gravity theories require the simultaneous satisfaction of: (i) the corresponding Einstein field equations, (ii) the matter field equations, and (iii) the Lorentz-violating field equations. Lorentz symmetry breaking can emerge spontaneously when tensor fields acquire non-zero vacuum expectation values. When frozen in these vacuum states, the dynamics of Lorentz-violating tensor fields may reduce to geometric constraints, potentially ruling out entire classes of compact objects. These constraints are crucial for ensuring physical consistency in Lorentz-violating frameworks, as they eliminate metric families incompatible with the preferred spacetime directions emerging from spontaneous Lorentz symmetry breaking. We investigate the criteria governing the emergence of these geometric constraints and analyze their consequences. Our analysis establishes a consistency framework for evaluating compact objects in these theories, demonstrating that several previously reported solutions in Lorentz-violating gravity models are physically inadmissible.
I Introduction
@@
General relativity (GR) remains the most successful theory for describing gravitational interaction at classical scales, grounded in the symmetries of diffeomorphism and local Lorentz invariance. However, the unification of GR with Quantum Mechanics still represents one of the greatest unsolved challenges in theoretical physics [1]. The energy scale relevant to quantum gravity—the Planck scale ($M_{\text{Pl}}=\sqrt{\hbar c/G}\approx 1.22\times 10^{19}\text{ GeV}/c^{2}$)—lies far beyond any experimentally accessible regime. Consequently, investigating quantum gravity necessitates alternative approaches. Theoretical models extending GR, including various approaches to Quantum Gravity, often predict that characteristic low-energy “relic signatures” should arise—among them, the violation of spacetime symmetries—, leading to observable deviations from standard GR predictions [2, 3, 4].
@@
These signatures, which might arise from quantum gravitational effects, could manifest as subtle deviations either in experiments of the Standard Model or GR [5, 6, 7, 8]. In this context, the possibility of Lorentz symmetry violation in an underlying theory—such as string theory [9, 10], loop quantum gravity [11, 12], noncommutative field theories [13] and Hořava gravity [14]—becomes particularly relevant, especially in the search for observational signatures of new physics. These questions can be investigated independently of specific models through effective field theory approaches [15, 16]. Thus, experimental tests of Lorentz violation in gravity not only validate extensions of GR but may also reveal evidence of physics beyond the current paradigm [17, 18, 19, 20].
@@
A promising approach to investigate Lorentz violation in pure gravity involves analyzing signatures induced in compact objects by preferred spacetime directions [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]—a phenomenon which can arise either spontaneously [36, 37] or explicitly [38, 39, 40]. Focusing on spontaneous Lorentz and diffeomorphism symmetry breaking, a potential term drives the dynamic evolution of a tensor field, leading to a nonzero vacuum expectation value (VEV). While being spontaneous, this violation preserves the Bianchi identities, despite implying both local Lorentz and diffeomorphism violation [16]. This compatibility with the geometric structure of GR makes spontaneous Lorentz and diffeomorphism symmetry breaking a natural candidate for exploring violation effects in strong gravity regimes, such as those found in neutron stars or black holes. Moreover, the spontaneous breaking of Lorentz symmetry leads to the emergence of Nambu-Goldstone modes and massive modes [41, 36]. The latter is generated through an alternative gravitational Higgs mechanism, assuming a smooth quadratic potential capable of producing new gravitational solutions [42, 43, 44].
@@
The aim of this work is to stress a point, sometimes missing in the discussion on compact objects in spontaneous Lorentz-violating gravity models, namely the self-consistency of compact objects in these frameworks. Crucially, self-consistent solutions of Lorentz-violating scenarios must simultaneously satisfy: (i) the corresponding Einstein field equations, (ii) the matter field equations, and (iii) the Lorentz-violating field equations. Remarkably, under certain conditions, the latter in a vacuum state can reduce to geometric constraints governing the metrics of compact objects. These constraints impose stringent requirements that rule out the emergence of families of compact objects and invalidate some previously reported solutions of Lorentz-violating gravity models.
@@
In the following, we analyze the emergence and implications of geometric constraints induced by spontaneous Lorentz and diffeomorphism violation in two prominent frameworks: (i) via a vector field (the bumblebee field) [16, 36], and (ii) through an antisymmetric rank-2 tensor field (sometimes known as the Kalb-Ramond field) [45]. The non-zero VEV of these fields introduces preferred spacetime directions, leading to small deviations in compact object solutions—including black holes [30, 31, 32, 33], neutron stars [34], and wormholes [35]—compared to standard GR predictions [^1] We note that compact object solutions have also been found in the metric-affine version of Einstein-bumblebee gravity [46, 47]. However, since we are working within the metric formulation, their analysis falls outside the scope of this work.. Our analysis plays a crucial role in the search for viable compact objects within Lorentz-violating gravity, as it allows us to probe various families of metric functions (including those mentioned above) and determine which are permitted by the Lorentz-violating fields.
The content of this paper is organized as follows. In Sec. II, we introduce the Lorentz-violating framework we will use, and derive its field equations. The assumptions governing Lorentz-violating field dynamics in the vacuum state for compact object geometries are presented in Sec. III. In Sec. IV, we investigate the self-consistent solutions for these compact objects generated by the VEV of the bumblebee field. In the following Sec. V, we perform the same analysis for the VEV of the antisymmetric rank-2 field. Finally, we summarize our results and discuss some perspectives in Sec. VI.
II Lorentz-violating framework
@@
The action that encodes Lorentz-violating effects from a vector field (bumblebee field) [16] and from an antisymmetric rank-2 tensor [45] in Riemann-spacetime limit can be written as
@@
where $\mathfrak{B}_{\mu\nu}=\partial_{\mu}\mathfrak{B}_{\nu}-\partial_{\nu}\mathfrak{B}_{\mu}$ and $H_{\lambda\mu\nu}=\partial_{\lambda}B_{\mu\nu}+\partial_{\mu}B_{\nu\lambda}+\partial_{\nu}B_{\lambda\mu}$ are, respectively, the field strength of the bumblebee and of the antisymmetric rank-2 tensor, $\tilde{V}(Y)$ and $V(X)$ are self-interaction potentials that drive the breaking of the Lorentz symmetry with $Y=\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}\pm\mathfrak{b}^{2}$, $X=B_{\mu\nu}B^{\mu\nu}\pm b^{2}$, where $\mathfrak{b}^{2}$ and $b^{2}$ are positive real constants (with dimension of length-2) and the $\pm$ sign denotes whether the nature of the vector $\mathfrak{B}^{\mu}$ and of the antisymmetric tensor ${B}_{\mu\nu}$ is timelike or spacelike. We assume that the self-interaction potentials $V$ and $\tilde{V}$ drive the formation of a nonzero vacuum value, respectively $\langle\mathfrak{B}_{\mu}\rangle=\mathfrak{b}_{\mu}$ and $\langle B_{\mu\nu}\rangle=b_{\mu\nu}$, which breaks local Lorentz and diffeomorphism symmetry [36, 37]. The vanishing VEVs $\langle V\rangle=0$ and $\langle\tilde{V}\rangle=0$ are ensured by the conditions $\mp\mathfrak{b}^{2}\equiv g^{\mu\nu}\mathfrak{b}_{\mu}\mathfrak{b}_{\nu}$ and $\mp b^{2}\equiv g^{\mu\alpha}g^{\nu\beta}b_{\mu\nu}b_{\alpha\beta}$, respectively, with $\langle\mathfrak{B}_{\mu}\mathfrak{B}^{\mu}\rangle=\mathfrak{b}_{\mu}\mathfrak{b}^{\mu}$ and $\langle B_{\mu\nu}B^{\mu\nu}\rangle=b_{\mu\nu}b^{\mu\nu}$.
Moreover, the additional terms in these Lagrangians—featuring non-minimal couplings between Lorentz-violating fields and spacetime curvature, with $R_{\alpha\beta\mu\nu}$ and $R_{\alpha\beta}$ denoting, respectively the Riemann and Ricci tensors and $\tilde{\xi}_{1,2}$ and ${\xi}_{1,2,3}$ are coupling constants (with dimension of length2)—represent the most widely studied models in the literature on Lorentz violation in the gravitational sector [36, 41, 45, 30, 31, 32, 33, 34, 35]. We recall that the bumblebee field does not couple to the Riemann tensor, as established in Ref. [16]. Moreover, we do not consider any coupling between the Lorentz-violating fields and the matter sector.
@@
where $T_{\mu\nu}=T_{\mu\nu}^{M}+T_{\mu\nu}^{LV}$ is the (total) energy-momentum tensor, with $T_{\mu\nu}^{M}$ and $T_{\mu\nu}^{LV}$ denoting the energy-momentum tensor of the matter fields and of the Lorentz-violating fields, respectively. We remark that $T_{\mu\nu}^{LV}$ is constructed by varying $S_{LV}=\int d^{4}x\sqrt{-g}(\mathcal{L}_{\mathfrak{B}}+\mathcal{L}_{{B}})$, containing, therefore, all the terms depending on $\mathfrak{B}_{\mu}$, $B_{\mu\nu}$ and their derivatives, along with all the non-minimal curvature-couplings controlled by $\tilde{\xi}_{1,2}$ and $\xi_{1,2,3}$. The non-zero VEVs of $\mathfrak{B}_{\mu}$ and $B_{\mu\nu}$ define a preferred direction in spacetime, which generates an anisotropic energy-momentum tensor [33], with different pressures along that direction compared to the orthogonal directions. Additionally, by varying the action (1) with respect to the matter fields (collectively denoted by $\psi_{m}$), one obtains the equations of motion of the matter fields. Independently, by varying the action (1) with respect to the Lorentz-violating fields, namely $\mathfrak{B}_{\mu}$ and $B_{\mu\nu}$, one obtains the field equations for the bumblebee field and for the antisymmetric rank-2 tensor field, respectively given by
$\nabla_{\alpha}\mathfrak{B}^{\alpha\nu}=$  $\,2\tilde{V}_{Y}\mathfrak{B}^{\nu}-\frac{\tilde{\xi}_{1}}{\kappa}R\mathfrak{B}^{\nu}-\frac{\tilde{\xi}_{2}}{\kappa}\mathfrak{B}_{\mu}R^{\mu\nu},$ (5)
$\nabla_{\alpha}H^{\alpha\mu\nu}=$  $\,4V_{X}B^{\mu\nu}-\frac{2\xi_{1}}{\kappa}RB^{\mu\nu}+\frac{2\xi_{2}}{\kappa}B_{\alpha}{}^{[\mu}R^{\nu]\alpha}$
$-\frac{2\xi_{3}}{\kappa}B_{\alpha\beta}R^{\alpha\beta\mu\nu}.$ (6)
@@
We emphasize that these consistency conditions can be understood through the lens of energy-momentum conservation. To ensure ordinary matter dynamics remain unaffected by Lorentz violation, we require separate conservation of the matter and Lorentz-violating energy-momentum tensors. Taking the covariant divergence of the Einstein equations and applying the contracted Bianchi identity $\nabla_{\mu}G^{\mu\nu}=0$ yields the on-shell condition:
$\nabla^{\mu}T_{\mu\nu}^{M}=0,\quad\nabla^{\mu}T_{\mu\nu}^{LV}=0.$ (7)
These conditions hold when the equations of motion for both matter and Lorentz-violating fields are independently satisfied, in conjunction with the appropriate Bianchi identities. Consequently, energy-momentum conservation is guaranteed [16, 45]. We emphasize that the energy-momentum tensor is automatically conserved in theories with spontaneous Lorentz and local diffeomorphism violation. A well-known consequence of this type of symmetry breaking is the emergence of massless Nambu-Goldstone (NG) modes, as well as the possible appearance of massive modes [41, 36]. Notably, this conservation persists even in hybrid theories that combine explicit diffeomorphism breaking with spontaneous Lorentz violation—but only under the condition that the Nambu-Goldstone mode remains dynamic while massive modes are suppressed [48, 49].
We note that if matter fields instead couple directly to Lorentz-violating fields, the individual conservation laws become inseparable due to their mutual dependence on the total energy-momentum tensor. In such scenarios, only the total conservation $\nabla^{\mu}T_{\mu\nu}=0$ can be enforced, and in this case one should modify the matter conservation.
III Compact objects in Lorentz-violating gravity
The search for astrophysical objects in Lorentz-violating scenarios is typically based on the assumption that the Lorentz-violating fields are frozen at their VEVs [30, 33, 31, 32]. Depending on the assumption about these VEVs, different families of compact objects may arise in scenarios with spontaneous local Lorentz and diffeomorphism symmetry breaking. It is important to point out that, imposing a particular form for the VEVs can turn the above dynamical equations into geometrical constraints, ruling out families of compact objects. These geometrical constraints emerge whenever the left-hand side of the Eqs. (5)-(II) vanish. This can happen either for vanishing or (covariant) divergence-free field strength configurations. The resulting equations, for non-vanishing VEVs, imposes constraints on the Ricci and Riemann tensor and on the Ricci scalar.
Notably, several compact objects reported as solutions of Lorentz-violating models are based on the former assumption, namely the field strengths of the bumblebee and of the antisymmetric rank-2 tensor, in the vacuum state,—respectively $\langle\mathfrak{B}_{\mu\nu}\rangle=\mathfrak{b}_{\mu\nu}$ and $\langle H_{\alpha\beta\gamma}\rangle=h_{\alpha\beta\gamma}$—, vanish [30, 33, 50, 51, 52, 53, 31, 32]. In the following we will show that this choice, namely
$\mathfrak{b}_{\mu\nu}=0,\quad h_{\alpha\beta\gamma}=0,$ (8)
can, instead of allowing, forbid the emergence of some of these compact objects. We will additionally see that geometric constraints are stringent, and, for particular choices of VEV states, only a few classes of metrics are allowed. Moreover, as our primary objective is to assess the validity of reported solutions, we intentionally adopt parameter choices—specifically, the coupling constants in the Lagrangian associated with the non-minimal couplings of the Lorentz-violating fields to gravity (e.g., couplings to the Ricci scalar, Ricci tensor, etc.)—consistent with those in prior studies seeking these solutions.